# Happy family of stable marriages

### 02 Apr 2026 v1

@@ 4.1 Invariance of the Self-Perimeter Under Affine Transformations
Assume the invariance holds for dimension n1n-1. The self-perimeter in $\mathbb{R}^{n}$ is recursively defined as:
$P(B^{(n)})=\int_{\partial^{*}B}\frac{V(B^{(n-1)}(\nu_{x}))}{\mathcal{H}_{n-1}(B^{(n-1)}(\nu_{x}))}d\mathcal{H}_{n-1}(x)$ (11)
where $\nu_{x}$ is the outward normal vector at $x\in\partial^{*}B$, and $B^{(n-1)}(\nu)=B\cap\Sigma(\nu)$ is the cross-section of $B$ on the hyperplane $\Sigma(\nu)$ passing through the origin and orthogonal to $\nu$.
Let $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be an invertible linear transformation, and let $\tilde{B}=T(B)$. For a point $x\in\partial^{*}B$, the corresponding point on the transformed boundary is $\tilde{x}=T(x)$. The normal vector at $\tilde{x}$ transforms according to the inverse transpose:
$\tilde{\nu}(\tilde{x})=\frac{(T^{-1})^{T}\nu_{x}}{\|(T^{-1})^{T}\nu_{x}\|}$ (12)
[…]
$\tilde{B}^{(n-1)}(\tilde{\nu})=\tilde{B}\cap\tilde{\Sigma}(\tilde{\nu})=T(B\cap\Sigma(\nu))=T(B^{(n-1)}(\nu))$ (13)
By our induction hypothesis, the self-volume in n1n-1 dimensions is affine-invariant, hence:
$V(\tilde{B}^{(n-1)}(\tilde{\nu}))=V(B^{(n-1)}(\nu))$ (14)
Next, we determine how the Euclidean measures transform. By Nanson’s formula (or the Jacobian of the surface transformation), the (n1)(n-1)-dimensional Hausdorff measure on the boundary transforms as:
$d\mathcal{H}_{n-1}(\tilde{x})=|\det(T)|\|(T^{-1})^{T}\nu_{x}\|d\mathcal{H}_{n-1}(x)$ (15)
Similarly, the Euclidean (n1)(n-1)-dimensional volume of the cross-section lying in the hyperplane $\Sigma(\nu)$ scales by the exact same geometric factor under the transformation $T$:
$\mathcal{H}_{n-1}(\tilde{B}^{(n-1)}(\tilde{\nu}))=|\det(T)|\|(T^{-1})^{T}\nu_{x}\|\mathcal{H}_{n-1}(B^{(n-1)}(\nu))$ (16)
Now, we evaluate the integral for the self-perimeter of the transformed body $\tilde{B}$:
[…]
$=\int_{\partial^{*}B}\frac{V(B^{(n-1)}(\nu_{x}))}{|\det(T)|\|(T^{-1})^{T}\nu_{x}\|\mathcal{H}_{n-1}(B^{(n-1)}(\nu_{x}))}\left(|\det(T)|\|(T^{-1})^{T}\nu_{x}\|d\mathcal{H}_{n-1}(x)\right)$ (17)
The Jacobian determinant and the norm of the transformed normal vector cancel out completely between the numerator and the denominator. We are left with:
$P(\tilde{B}^{(n)})=\int_{\partial^{*}B}\frac{V(B^{(n-1)}(\nu_{x}))}{\mathcal{H}_{n-1}(B^{(n-1)}(\nu_{x}))}d\mathcal{H}_{n-1}(x)=P(B^{(n)})$ (18)
Consequently, since $V(B^{(n)})=\frac{1}{n}P(B^{(n)})$, the self-volume is also invariant under affine transformations preserving the origin. This completes the proof. ∎
4.2 Examples
The hypercube
The central cross-section of $B_{n}$ passing through the origin and orthogonal to this normal is $B_{n-1}=[-1,1]^{n-1}$. The Euclidean (n1)(n-1)-dimensional volume of this cross-section is $\mathcal{H}_{n-1}(B_{n-1})=2^{n-1}$.
Because the normal vector is constant across a single face, the integrand is also constant. The integral over the boundary $\partial B_{n}$ is the sum of the integrals over the $2n$ faces:
$P(B_{n})=\sum_{k=1}^{2n}\int_{\text{face}_{k}}\frac{V(B_{n-1})}{2^{n-1}}d\mathcal{H}_{n-1}(x)$
The Euclidean (n1)(n-1)-dimensional volume of each face is also $2^{n-1}$. Thus, the integral evaluates to:
$P(B_{n})=\sum_{k=1}^{2n}\left(\frac{V(B_{n-1})}{2^{n-1}}\times 2^{n-1}\right)=2nV(B_{n-1})$
The self-volume is postulated as $V_{n}=P_{n}/n$. Substituting the perimeter yields:
$V(B_{n})=\frac{2nV(B_{n-1})}{n}=2V(B_{n-1})$
With the base case for the one-dimensional segment being $V(B^{(1)})=2$, solving the recurrence gives:
$V([-1,1]^{n})=2^{n}$
It is interesting to compare the perimeter to volume ration according to Busemann’s definition. The Busemann volume for the Minkowski unit ball $B=[-1,1]^{n}$ is $\omega_{n}$.
The corresponding surface area is then obtained by integrating the Euclidean surface element, normalized by the volume of the central cross-section of the unit ball parallel to the tangent plane.
For the hypercube $B=[-1,1]^{n}$:
  • The boundary consists of $2n$ faces.
  • For each face, the central cross-section parallel to it is an (n1)(n-1)-dimensional hypercube $1,1]^{n-1}$.
  • The Euclidean (n1)(n-1)-dimensional volume of this central cross-section is $2^{n-1}$.
  • The normalization factor applied to the Euclidean surface element $d\mathcal{H}_{n-1}$ is $\frac{\omega_{n-1}}{2^{n-1}}$.
Integrating this constant factor over a single face (which has a Euclidean area of $2^{n-1}$):
$\int_{\text{face}}\frac{\omega_{n-1}}{2^{n-1}}d\mathcal{H}_{n-1}(x)=\frac{\omega_{n-1}}{2^{n-1}}\times 2^{n-1}=\omega_{n-1}$
Summing across all $2n$ faces provides the total Busemann surface area:
$P_{Bus}(B)=2n\omega_{n-1}$
The ratio of the Busemann surface area to the Busemann volume for the hypercube is:
$\frac{P_{Bus}(B)}{V_{Bus}(B)}=\frac{2n\omega_{n-1}}{\omega_{n}}$
compared to $P(B)/V(B)=n$ according to our definition. Note that, as $n\rightarrow\infty$, $\omega_{n-1}/\omega_{n}\approx\sqrt{n/2\pi}$ so $P_{Bus}/V_{Bus}\approx\sqrt{2/\pi}n^{3/2}$.
The Euclidean ball
Since the Lebesgue measure of the cross sections of the Euclidean ball are all constants equal to $\omega_{n-1}$, it follows that the self perimeter is
$P(B_{E}^{n})=\int_{{\mathbb{S}^{n-1}}}\frac{V(B_{R}^{n-1})}{\omega_{n-1}}d{\cal H}_{n-1}(x)=\frac{V(B_{R}^{n-1})}{\omega_{n-1}}\int_{{\mathbb{S}^{n-1}}}d{\cal H}_{n-1}(x)\ .$
The last integral is just the perimeter of the sphere so is equal $n\omega_{n}$. Thus $P(B_{E}^{n})=nV(B_{E}^{n-1})\frac{\omega_{n}}{\omega_{n-1}}$. By our definition $P(B_{E}^{n})=nV(B_{E}^{n})$ so
$V(B_{E}^{n})=V(B_{R}^{n-1})\ .\frac{\omega_{n}}{\omega_{n-1}}$
As we know $V(B_{E}^{2})=\omega_{2}$ so we get the agreement $V(B_{E}^{n})=\omega_{n}$ for any $n\geq 2$ between our definition and the Lebesgue measure of the Euclidean ball.
Open Questions
  • What is the sup/inf values of $V_{B}^{(n)}$ for $n>2$ among all CCS $B\in\mathbb{R}^{n}$?
  • Is there a CCS set $B\subset\mathbb{R}^{n}$ which minimize (maximize) $V_{B}^{(n)}$?
5 Uniqueness of optimal center
5.1 Convexity of the Self-Perimeter Functional for an $n$-Simplex
In this section, we extend the observation made in $\mathbb{R}^{2}$ to an arbitrary $n$-dimensional simplex $\Delta^{n}$, proving that the self-perimeter functional $P(p)$ derived from Equation (11) is strictly convex with respect to the interior reference point $p$. This guarantees the existence of a unique minimizing point within the simplex.
Let $\Delta^{n}$ be an $n$-dimensional simplex with $n+1$ vertices. Any interior point $p\in\text{int}(\Delta^{n})$ can be uniquely represented by its strictly positive barycentric coordinates $\lambda=(\lambda_{1},\dots,\lambda_{n+1})$, where $\sum_{i=1}^{n+1}\lambda_{i}=1$ and $\lambda_{i}>0$ for all $i$. Let $F_{i}$ denote the (n1)(n-1)-dimensional face opposite to the $i$-th vertex.
By applying the recursive definition of the self-perimeter (11), the integral over the boundary $\partial\Delta^{n}$ decomposes into a sum over the $n+1$ faces. For a specific face $F_{i}$, the integration is evaluated over the (n1)(n-1)-dimensional cross-section $S_{i}$ passing through $p$ and parallel to $F_{i}$.
Because $S_{i}$ is parallel to $F_{i}$, it is a simplex similar to $F_{i}$ with a linear scaling factor of $(1-\lambda_{i})$. Consequently, the ratio of their Euclidean volumes, which appears in the denominator of Equation (11), is given by:
To evaluate the self-volume of the cross-section $S_{i}$ (the numerator in Equation 11), we first determine the barycentric coordinates of $p$ within $S_{i}$. By redistributing the weights among the remaining $n$ vertices, the local coordinates are $\mu_{j}=\frac{\lambda_{j}}{1-\lambda_{i}}$ for $j\neq i$. By the induction hypothesis, the self-perimeter of the (n1)(n-1)-simplex $S_{i}$ is proportional to the sum of the reciprocals of its barycentric coordinates:
$V^{(n-1)}(S_{i})=\sum_{j\neq i}\frac{1}{\mu_{j}}=(1-\lambda_{i})\sum_{j\neq i}\frac{1}{\lambda_{j}}$ (20)
Substituting these components back into the sum over all faces yields the full functional for the $n$-simplex:
$P(\Delta^{n}-p)=\sum_{i=1}^{n+1}\left[\frac{1}{(1-\lambda_{i})^{n-1}}\cdot(1-\lambda_{i})\sum_{j\neq i}\frac{1}{\lambda_{j}}\right]=\sum_{i=1}^{n+1}\sum_{j\neq i}\frac{1}{(1-\lambda_{i})^{n-2}\lambda_{j}}$ (21)
Notice that all Euclidean length and area metrics cancel out, leaving a functional that depends exclusively on the affine geometric parameters (the barycentric coordinates).
Note that Busemann definition [2] (c.f. (1) ) of the perimeter of the simplex is:
$P_{\text{Busemann}}(\Delta^{n}-p)=\omega_{n-1}\sum_{i=1}^{n+1}\frac{1}{(1-\lambda_{i})^{n-1}}$ (22)
5.1.1 Proof of Strict Convexity
To prove that $f(p):=P(\Delta^{n}-p)$ is strictly convex, we analyze the individual terms in the double sum. Let $h_{ij}(\lambda)=\frac{1}{(1-\lambda_{i})^{n-2}\lambda_{j}}$. Since the sum of strictly convex functions is strictly convex, it suffices to observe that each $h_{ij}$ is strictly convex in the variables $\lambda_{i},\lambda_{j}$.
Consequently, $P(\Delta^{n}-p)$, being a sum of strictly convex functions, is itself strictly convex on the domain of valid barycentric coordinates. This completes the proof and guarantees that the self-perimeter functional possesses a unique minimum inside the $n$-simplex.
5.2 Strict Convexity for Cartesian Products
@@
$V^{(n+m)}(A\times B)=V^{(n)}(A)V^{(m)}(B)$ (23)
From this, Theorem 3.2 and Section 5.1 we obtain:
Corollary 5.1.
If $A\in\mathbb{R}^{n}$ is a Cartesian product of convex sets of dimension 2, intervals of dimension 1 and simplexes of any dimension, then $p\rightarrow P(A-\{p\})$ is a strongly convex set in $Int(A)$. In particular, there exists a unique $p\in Int(A)$ which minimize $P(A-\{p\})$.
@@
Let $n=1$ and $m=1$. Both $A$ and $B$ are one-dimensional segments, denoted $I_{1}$ and $I_{2}$. The Cartesian product $I_{1}\times I_{2}$ is a 2D rectangle. The boundary $\partial(I_{1}\times I_{2})$ decomposes into two parts: $(\partial I_{1}\times I_{2})\cup(I_{1}\times\partial I_{2})$.
For the faces $\partial I_{1}\times I_{2}$, the normal vectors are parallel to $I_{1}$. The orthogonal cross-section is $I_{2}$. The Euclidean length of these two faces is $2|I_{2}|$. The contribution to the self-perimeter is:
$\int_{\partial I_{1}\times I_{2}}\frac{V^{(1)}(I_{2})}{|I_{2}|}d\mathcal{H}_{1}=2|I_{2}|\frac{V^{(1)}(I_{2})}{|I_{2}|}=2V^{(1)}(I_{2})$
By symmetry, the integral over $I_{1}\times\partial I_{2}$ evaluates to $2V^{(1)}(I_{1})$. Thus, the total self-perimeter is:
$P^{(2)}(I_{1}\times I_{2})=2V^{(1)}(I_{2})+2V^{(1)}(I_{1})$
Since $V^{(1)}(I)=2$ for any segment, $P^{(2)}(I_{1}\times I_{2})=2(2)+2(2)=8$. The self-volume is $V^{(2)}=P^{(2)}/2=4$. Since $V^{(1)}(I_{1})V^{(1)}(I_{2})=2\times 2=4$, the base case holds.
@@
We evaluate the self-perimeter integral over these two components separately.
Part 1: Integration over $\partial A\times B$
For any point $(x,y)\in\partial A\times B$, the outward normal vector $\nu_{x}$ lies entirely within the subspace $\mathbb{R}^{n}$. The orthogonal cross-section is given by $A_{x}\times B$, where $A_{x}=A\cap\Sigma(\nu_{x})$ is an (n1)(n-1)-dimensional convex set. The sum of the dimensions of $A_{x}$ and $B$ is (n1)+m=k1(n-1)+m=k-1. By the inductive hypothesis, we can decompose its volume:
$V^{(k-1)}(A_{x}\times B)=V^{(n-1)}(A_{x})V^{(m)}(B)$
@@
$I_{1}$ $=\int_{\partial A\times B}\frac{V^{(k-1)}(A_{x}\times B)}{\mathcal{H}_{k-1}(A_{x}\times B)}d\mathcal{H}_{k-1}$
$=\int_{\partial A}\int_{B}\frac{V^{(n-1)}(A_{x})V^{(m)}(B)}{|A_{x}||B|}dy\,d\mathcal{H}_{n-1}(x)$
@@
$I_{1}$ $=V^{(m)}(B)\int_{\partial A}\frac{V^{(n-1)}(A_{x})}{|A_{x}|}d\mathcal{H}_{n-1}(x)$
By the recursive definition, the remaining integral is exactly the self-perimeter of $A$, denoted $P^{(n)}(A)$. Since $P^{(n)}(A)=nV^{(n)}(A)$, we obtain:
$I_{1}=V^{(m)}(B)P^{(n)}(A)=nV^{(n)}(A)V^{(m)}(B)$
Part 2: Integration over $A\times\partial B$
By symmetry, for points $(x,y)\in A\times\partial B$, the normal vector lies entirely in $\mathbb{R}^{m}$. The cross-section is $A\times B_{y}$, where $B_{y}$ is an (m1)(m-1)-dimensional slice of $B$. Applying the inductive hypothesis again (since n+(m1)=k1n+(m-1)=k-1):
$V^{(k-1)}(A\times B_{y})=V^{(n)}(A)V^{(m-1)}(B_{y})$
@@
$I_{2}$ $=\int_{A\times\partial B}\frac{V^{(n)}(A)V^{(m-1)}(B_{y})}{|A||B_{y}|}dx\,d\mathcal{H}_{m-1}(y)$
$=V^{(n)}(A)\int_{\partial B}\frac{V^{(m-1)}(B_{y})}{|B_{y}|}d\mathcal{H}_{m-1}(y)$
$=V^{(n)}(A)P^{(m)}(B)$
$=mV^{(n)}(A)V^{(m)}(B)$
Conclusion of the Inductive Step:
Summing the two contributions gives the total self-perimeter of the Cartesian product:
$P^{(n+m)}(A\times B)$ $=I_{1}+I_{2}$
$=nV^{(n)}(A)V^{(m)}(B)+mV^{(n)}(A)V^{(m)}(B)$
$=(n+m)V^{(n)}(A)V^{(m)}(B)$
Finally, using the postulated relationship between self-perimeter and self-volume, we divide by the combined dimension $n+m$:
$V^{(n+m)}(A\times B)$ $=\frac{P^{(n+m)}(A\times B)}{n+m}$
$=\frac{(n+m)V^{(n)}(A)V^{(m)}(B)}{n+m}$
$=V^{(n)}(A)V^{(m)}(B)$
This completes the induction step. The theorem holds for all dimensions $n,m\geq 1$.
6 Alexandrov’s type problem for self-surface area

### 09 Apr 2026 v2

@@ 4.2 Convexity of self-volume
Strict convexity follows from the log-convexity of $\log(f)$ where $f(\vec{\lambda})=\prod_{m=1}^{n-1}\frac{1}{1-S_{m}}$ and $S_{m}=\sum_{i=1}^{m}\lambda_{k_{i}}$. Hence, the minimum is unique. Since this function is invariant under permutation of $\lambda_{i}$, the minimum must attain where all $\lambda_{i}$ are equal, namely $\lambda_{i}=(n+1)^{-1}$. ∎
@@
$\omega^{(n+m)}_{A\times B}=\omega^{(n)}_{A}\omega^{(m)}_{B}$ (15)
@@
Let $n=1$ and $m=1$. Both AA and $B$ are one-dimensional segments, denoted $I_{1}$ and $I_{2}$. The Cartesian product $I_{1}\times I_{2}$ is a 2D rectangle. The boundary $\partial(I_{1}\times I_{2})$ decomposes into two parts: $(\partial I_{1}\times I_{2})\cup(I_{1}\times\partial I_{2})$. For the faces $\partial I_{1}\times I_{2}$, the normal vectors are parallel to $I_{1}$. The orthogonal cross-section is $I_{2}$. The Euclidean length of these two faces is $2|I_{2}|$. The contribution to the self-perimeter is:
$\int_{\partial I_{1}\times I_{2}}\frac{\omega^{(1)}_{I_{2}}}{|I_{2}|}d\mathcal{H}_{1}=2|I_{2}|\frac{\omega^{(1)}_{I_{2}}}{|I_{2}|}=2\omega^{(1)}_{I_{2}}$
By symmetry, the integral over $I_{1}\times\partial I_{2}$ evaluates to $2\omega^{(1)}_{I_{1}}$. Thus, the total self-perimeter is:
$\omega^{(2)}_{I_{1}\times I_{2}}=\omega^{(1)}_{I_{2}}+\omega^{(1)}_{I_{1}}=4$
This is consistent with the self-volume of $1,1]^{2}$ via Section 4.1.
@@
We evaluate the self-perimeter integral over these two components separately. For any point $(x,y)\in\partial A\times B$, the outward normal vector $\nu_{x}$ lies entirely within the subspace $\mathbb{R}^{n}$. The orthogonal cross-section is given by $A_{x}\times B$, where $A_{x}=A\cap\Sigma(\nu_{x})$ is an (n1)(n-1)-dimensional convex set. The sum of the dimensions of $A_{x}$ and $B$ is (n1)+m=k1(n-1)+m=k-1. By the inductive hypothesis, we can decompose its volume:
$\omega^{(k-1)}_{A_{x}\times B}=\omega^{(n-1)}_{A_{x}}\omega^{(m)}_{B}$
@@
$I_{1}$ $=\int_{\partial A\times B}\frac{\omega^{(k-1)}_{A_{x}\times B}}{\mathcal{H}_{k-1}(A_{x}\times B)}d\mathcal{H}_{k-1}$
$=\int_{\partial A}\int_{B}\frac{\omega^{(n-1)}_{A_{x}}\omega^{(m)}_{B}}{|A_{x}||B|}dy\,d\mathcal{H}_{n-1}(x)$
@@
$I_{1}$ $=\omega^{(m)}_{B}\int_{\partial A}\frac{\omega^{(n-1)}_{A_{x}}}{|A_{x}|}d\mathcal{H}_{n-1}(x)$
By the recursive definition, the remaining integral is exactly the self-perimeter of AA, denoted $P^{(n)}(A)$. Since $P^{(n)}(A)=n\omega^{(n)}_{A}$, we obtain:
$I_{1}=n\omega^{(n)}_{A}\omega^{(m)}_{B}$
Similarly:
$\omega^{(k-1)}_{A\times B_{y}}=\omega^{(n)}_{A}\omega^{(m-1)}_{B_{y}}$
@@
$I_{2}=\int_{A\times\partial B}\frac{\omega^{(n)}_{A}\omega^{(m-1)}_{B_{y}}}{|A||B_{y}|}dx\,d\mathcal{H}_{m-1}(y)=m\omega^{(n)}_{A}\omega^{(m)}_{B}$
So:
$\omega^{(n+m)}_{A\times B}=(m+n)^{-1}(I_{1}+I_{2})=\omega^{(n)}_{A}\omega^{(m)}_{B}\$.
4.3 Strict Convexity for Cartesian Products
From Theorem 4.2 and Theorem 3.3 we obtain:
Corollary 4.2.
If $A\in\mathbb{R}^{n}$ is a Cartesian product of convex sets of dimension 2, intervals of dimension 1, and simplices of any dimension, then $p\rightarrow\omega^{(n)}(A-\{p\})$ is a strongly convex function in $Int(A)$. In particular, there exists a unique $p\in Int(A)$ which minimizes $\omega^{(n)}(A-\{p\})$.
4.4 Conjecture
  • The CCS (convex centrally symmetric) body in $\mathbb{R}^{n}$ of maximal self-volume is the $n$-cube $\omega^{(n)}_{B}=2^{n}$.
  • The non-centrally symmetric, convex body in $\mathbb{R}^{n}$ of maximal self-volume with respect to its centroid is the $n$-simplex $B=\Delta^{n}$ ($\omega^{(n)}_{B}=\frac{(n+1)^{n}}{n!}$).
  • If $n=2k$, then the CCS body of minimal self-volume is the Cartesian product of $n$ hexagons ($\omega^{(n)}_{B}=3^{k}$). If $n=2k+1$, it is the Cartesian product of $k$ hexagons and a 1D interval ($\omega^{(n)}_{B}=2\times 3^{k}$).
4.5 Invariance of the Self-Perimeter Under Affine Transformations
Let $K\subset\mathbb{R}^{n}$ be a convex body with the origin in its interior. For a point $x\in\partial K$, let $\tau_{x}$ denote the tangent hyperplane at $x$, and let $K\cap\tau_{x}^{\perp}$ denote the central cross-section parallel to $\tau_{x}$. Our recursive formula defines the $n$-dimensional self-perimeter, $P_{n}(K)$, by integrating the (n1)(n-1)-dimensional self-perimeter of the parallel central cross-sections over the boundary, weighted by the Busemann surface element:
$P_{n}(K)=\int_{\partial K}P_{n-1}(K\cap\tau_{x}^{\perp})\frac{d\mathcal{H}^{n-1}(x)}{\text{Vol}_{n-1}(K\cap\tau_{x}^{\perp})}$ (16)
For the base case $n=1$, the body $K$ is a line segment, and its 1-dimensional self-perimeter is defined as a constant $P_{1}(K)=c$ (typically $c=2$).
Proof of Affine Invariance
We rely on the following geometric property established by Busemann, which we quote without proof:
Theorem 4.3.
[Busemann] The ratio of the Euclidean surface measure to the volume of the parallel central cross-section is invariant under any non-singular linear transformation $A:\mathbb{R}^{k}\to\mathbb{R}^{k}$. Specifically, both the surface element $d\mathcal{H}^{k-1}(x)$ and the cross-sectional volume ${\cal H}_{k-1}(K\cap\tau_{x}^{\perp})$ scale by the identical Jacobian factor $J=|\det A|\,|A^{-t}n_{x}|$, leading to exact cancellation:
$\frac{d\mathcal{H}_{k-1}(Ax)}{{\cal H}_{k-1}(A(K)\cap A(\tau_{x})^{\perp})}=\frac{d\mathcal{H}^{k-1}(x)}{{\cal H}_{k-1}(K\cap\tau_{x}^{\perp})}$ (17)
Theorem $4.$
For any $n\geq 1$ and any non-singular affine transformation $T(x)=Ax+b$, the recursive self-volume $\omega^{(n)}_{B}$ is invariant: $\omega^{(n)}(B)=\omega^{(n)}(T(B))$.
The proof of Theorem 4.4 follows from Busemann’s Theorem 4.3 and an induction argument based on the recursive Definition 4.1.
5 Alexandrov-type problem for self-surface area